I must solve the system of equations by graphing.
x-y=-2
x+y=12
I cancelled out the y's so what was left was 2x=8- then I divide by 2 so x=4. Now how do I get y, and how do I figure out the ordered pair? This is only one number. This is where I get confused, and have a problem figuring out how to graph it.

Answer by Alan3354(69443)

x-y=-2
x+y=12
I cancelled out the y's so what was left was 2x=8- then I divide by 2 so x=4. Now how do I get y, and how do I figure out the ordered pair? This is only one number. This is where I get confused, and have a problem figuring out how to graph it.
-------------
x-y=-2
x+y=12
------ Add the 2 eqns
2x = 10
x = 5
-----
y = 7
---> (5,7)
--------------
It says "solve by graphing", tho.
dl the FREE software at www.padowan.dk.com/graph/
graph the 2 eqns, see where they intersect.